OpenStudy (chris215):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 2, -4, and 1 + 3i I know the polynomial is (x-2)(x + 4)(x - (1+3i) but do I do next?

OpenStudy (thomas5267):

Note that is not the whole list of zeros. This is because for a real coefficient polynomial, complex roots must come in conjugates.

OpenStudy (chris215):

ok but im not sure what to do next

OpenStudy (welshfella):

do you know what the conjugate of 1 + 3i is?

OpenStudy (chris215):

no

OpenStudy (welshfella):

just replace the + by -

OpenStudy (welshfella):

1 - 3i

OpenStudy (chris215):

ok so I got f(x) = x^4 - 3x^3 + 6x^2 - 18x + 80

OpenStudy (thomas5267):

I don't think the answer is correct. Try again and remember that \((1-3i)(1+3i)=1^2-(3i)^2=1-(-9)=10\).

OpenStudy (welshfella):

( x - (1+ 3i) (x - (1-3i) = x^2 -x(1 + 3i) - x(1 - 3i0 + (1+31)(1-31) = x^2 - 2x + 10